OpenStudy (anonymous):

Given f(x), complete the equation for the inverse of f(x). F(x)=6x-7 , f^-1(y)=?

OpenStudy (anonymous):

solve \[y=6x-7\] for \(x\) in two steps \[y+7=6x\] \[\frac{y+7}{6}=x\] so \(f^{-1}(y)=\frac{y+7}{6}\)